OpenStudy (anonymous):

please help, I will give medal and fan who ever helps me 2/3 * x = 4/5 Multiply both sides by reciprocal of 2/3, which is 3/2 3/2 * 2/3 * x = 3/2 * 4/5 Simplify: x = 3/1 * 2/5 Multiply: x = 6/5 = 1 ¹/₅ −4d/3 = 3/25 Multiply both sides by 3 −4d = 9/25 Divide both sides by coefficient of d (−4) d = −9/100 −8/9 * c = −24 c = 24 * 9/8 c = 3 * 9/1 c = 27 3/4 * a = 12 a = 12 * 4/3 a = 16

OpenStudy (anonymous):

Explain each step. You must use the terms reciprocal and coefficient at least once. Label your steps with Step 1, Step 2, Step 3, and Step 4.

OpenStudy (anonymous):

2/3 * x = 4/5 Step 1: Multiply both sides with 3/2 which is the reciprocal of 2/3, which is the coefficient of the x 3/2 * 2/3 * x = 3/2 * 4/5 Step 2: Simplify: x = 3/1 * 2/5 Step 3: Multiply: x = 6/5 = 1 ¹/₅ −4d/3 = 3/25 Step 1: Multiply both sides by -3/4 which is the reciprocal of (-4/3), which is the coefficient of d d = 9/25 *(-3/4) d = −27/100 −8/9 * c = −24 Step 1: Multiply both sides with -9/8 which is the reciprocal of (-8/9), which is the coefficient of c c = 24 * 9/8 Step 2: Simplify c = 3 * 9/1 Step 3: Multiply c = 27 3/4 * a = 12 Step 1: Multiply both sides with 4/3 which is the reciprocal of 3/4, which is the coefficient of a a = 12 * 4/3 step 2: simplify a = 16

OpenStudy (anonymous):

step 4 is there, right?

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

Is it compulsory to have step 4?

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

I mainly need the justification for them

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

Well you have the justification

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Every step is justified. You just dont have the number of steps.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

anytime.